OpenStudy (anonymous):

y varies jointly with p and q. when p=2 and q=3 and y=48. find y when p=3 and q=4

OpenStudy (anonymous):

@viktor.ray the anwsers given are y=12 , y=24, y=48, and y=96

OpenStudy (anonymous):

oops misunderstood the question slightly y=k*p*q 48=k*2*3-->k=8 y=8*3*4=96 (Answer)

OpenStudy (anonymous):

haha its all good (: thank youu!